# Investigating the transverse-momentum- and pseudorapidity-dependent flow vector decorrelation in p–Pb collisions with a Multi-Phase Transport model

### v1

@@ IV.1 Transverse-momentum dependence of decorrelation
We first examine the $p_{\mathrm{T}}$-differential $v_{2}$ of charged hadrons in the pseudorapidity region $|\eta|<$ 0.8 using the AMPT model with the default configuration (Par1) as described in Table 1. The $v_{2}\{2\}$ and $v_{2}[2]$ are calculated according to Eq. 4 and Eq. 5. Figure 1 shows the $p_{\mathrm{T}}$-differential $v_{2}\{2\}$ and $v_{2}[2\}$ for 0–20%, 20–40%, and 40–60% centrality bins obtained from the AMPT calculations, compared with ALICE measurements [41]. To suppress contributions from short-range jet correlations, the $|\Delta\eta|>0.8$ is applied. The calculations reasonably reproduce the $v_{2}$ data from central to semi-central collisions, and the ratio $v_{2}\{2\}/v_{2}[2\}$ shows deviations from unity above $\it{p}_{\mathrm{T}}\simeq$ 2 GeV/$c$. These deviations increase with $\it{p}_{\mathrm{T}}$, similar to findings in Pb–Pb collisions, indicating $p_{\mathrm{T}}$-dependent flow vector fluctuations in p–Pb collisions from the AMPT model. Unlike Pb–Pb collisions, these deviations do not significantly depend on centrality selections. To determine if these deviations are caused by non-flow effects from long-range jet correlations, we apply the template fit subtraction method described in Eq. 10. As expected, the $v_{2}$ after subtraction, shown in Fig. 1 (dashed line), is lower than the results without subtraction, especially at high $p_{\mathrm{T}}>$ 2 GeV/$c$. The ratio $v_{2}\{2\}/v_{2}[2\}$ after subtraction is consistent with that before subtraction within uncertainties. This indicates that the long-range jet contribution to the decorrelations is negligible here, and the observed $p_{\mathrm{T}}$-dependent flow vector fluctuations in p–Pb collisions are mainly from real flow signals.
Figure 2 shows the $p_{\mathrm{T}}$-dependent $v_{2}\{2\}/v_{2}[2\}$ ratio in 0–20%, 20–40%, and 40–60% centrality classes obtained from the AMPT model with different configurations, compared to the ALICE data points [41]. For direct comparison with the data, nonflow subtraction is not applied here. The impact of the partonic scattering on the ratio can be seen by comparing the results from Par1, Par2, and Par3, as shown in Fig. 2 (top). The ratio obtained from larger parton scattering cross section (Par2) is slightly higher than the results from Par1. On the other hand, a large deviation to unity is observed when the cross section is set to $\sigma=0$ mb (Par3). It may attribute to the break of factorization from the nonflow, since the parton scattering process was proved to be the main contribution to the flow in small collision systems [70]. The effect from varying the initial conditions via the Lund string parameters, $a$ and $b$, can be seen by comparing the results from Par4 and Par1, as shown in Fig. 2 (top). One can see that the $v_{2}\{2\}$/$v_{2}[2]$ ratio obtained with larger Lund $a$ and $b$ (Par4) is lower in the most 0–20% collisions. With the increaing of the centralities, the resuts from Par4 are enhanced, and finally are hihger than the results from Par1. The effects from the hadronic scatterings are studied by varying the cutoff time of hadronic rescatterings, which can seen in the comparison of the results of Par1, Par5 and Par6, as shown in Fig. 2 (bottom). The $v_{2}\{2\}$/$v_{2}[2]$ ratio from Par1 and Par6 are consistent within uncertainties from the central to peripheral collisions, indicating that the impact from hadronic scatterings on the flow vector fluctuations are negligible. It is consistent with the previous findings that the hadronic scatterings have almost no effects on the $p_{\mathrm{T}}$-differential $v_{2}$ in p–Pb collisions [70]. For the case of Par5, where both the resonance decay and the hadronic scatterings are turn off, the $v_{2}\{2\}$/$v_{2}[2]$ ratio is lower at $p_{\mathrm{T}}<0.8$ GeV/$c$ but higher at $p_{\mathrm{T}}>1.4$ GeV/$c$ compare to the results of Par1 and Par6. It is expected since the decay processes convert high-$p_{\mathrm{T}}$ hadrons into low-$p_{\mathrm{T}}$ particles, thereby shifting the overall ratio toward the low-$p_{\mathrm{T}}$ region. One can see that the results of all configurations underestimate the data, suggesting that the model needs further tuning.
The factorization ratio $r_{2}$, definded in Eq. 7, is calculated in different centrality and $p^{b}_{\mathrm{T}}$ ranges, and compared to the ALICE data points [41] in Fig. 3. The results from the AMPT model with different configurations are shown in Fig. 3(a) (Par1, Par2 vs, Par3) and Fig. 4(a) (Par1, Par4, Par5, Par6). Simlar to the Fig. 2, nonflow subtraction is not applied here. The model with a 0.5 mb cross section, $a$ = 0.3, $b$ = 0.15 and $\tau$ = 30 fm/$c$ (Par1), provides a fair description of the data in all centrality and $p_{\mathrm{T}}$ bins. The deviation of the $r_{2}$ from the unity has no dependence on the centrality selections, and is more pronounced with the increasing $|p^{a}_{\mathrm{T}}-p^{b}_{\mathrm{T}}|$. The $r_{2}$ obtained with larger Lund $a$ and $b$ (Par4) are higher in more peripheral collisions, reflecting the effects from the initial conditions. Changing the partonic scattering cross section from 0.5 mb to 3 mb (Par1 vs Par3) slightly enhance the $r_{2}$, while the results obtained with $\sigma=0$ mb shows a significant breakdown of the factorization at large $|p^{a}_{\mathrm{T}}-p^{b}_{\mathrm{T}}|$. The impacts from the hadronic scatterings are negligible for all centralities and $p_{\mathrm{T}}$ bins (Par1 vs Par6), but the resonance decay process still has large impacts on the $r_{2}$ at low-$p_{\mathrm{T}}$ region (Par1 vs Par5). One can see that the calculations of $r_{2}$ and $v_{2}\{2\}$/$v_{2}[2]$ ratio have generally similar dependence on the initial conditions, parton scattering cross section, hadron scattering process. It is expected since $r_{2}$ is basically the double-differential $v_{2}\{2\}$/$v_{2}[2]$ ratio [59].
(a) The AMPT calculations are performed with different cross setions $\sigma$ of parton scatterings and Lund parameters $a$ and $b$.
@@
Figure 5: (Color online) The $v_{2}$ as a function of pseudorapidity $\eta$ in in different centrality ranges, obtained from the AMPT model with various configurations. ALICE data points [54] are shown as black open circles for comparison.
(a) The AMPT calculations are performed with different cutoff time $\tau$.
Figure 6: (Color online) The $v_{2}$ as a function of pseudorapidity $\eta$ in in different centrality ranges, obtained from the AMPT model with various configurations. ALICE data points [54] are shown as black open circles for comparison.
@@
Figure 7: (Color online) The square root of the product of factorization ratios, $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$, as a function of $\eta^{a}$ for 3.0 $<\eta^{b}<$ 4.0 (left) and $5.04.4<\eta^{b}<$ 5.0 (right), in 0–20% centrality class of p–Pb collisions at 5.02 TeV, obtained from the AMPT model with various tunings. The CMS data points [42] are shown as black open circles for comparison.
@@ IV.2 Pseudorapidity-dependence of decorrelation
The $\eta$-dependent $v_{2}$ is first calculated before studying the decorrelation. We follow the two-particle correlation method and the nonflow subtraction strategy used by the ALICE experiment [54]. The correlation between three groups of particles is constructed based on Eq. 10, then the flow coefficients are extracted for three combinations:
$v_{n}(\eta^{a})=\sqrt{\frac{V_{n\Delta}(\eta^{a},\eta^{b})V_{n\Delta}(\eta^{a},\eta^{c})}{V_{n\Delta}(\eta^{b},\eta^{c})}},$ (11)
@@
where $a$, $b$, and $c$ represent the hadrons with $p_{\mathrm{T}}>$ 0 GeV/$c$ at different rapidity bins. Similarly, $v_{n}(\eta^{a})$ and $v_{n}(\eta^{b})$ can be obtained by alternating the indices $a$, $b$, and $c$ in Eq. 11. Figure 5 shows the $v_{2}$ of the charged hadrons for $p_{\mathrm{T}}>$ 0 GeV/$c$ as a function of $\eta$ in 0–5%, 5–10%, 10–20%, and 20–40% centrality ranges, and the comparison to the ALICE data [54]. The results obtained from different configurations of the AMPT model are shown in Fig. 5(a) and Fig. 6(a), respectively. The calculations obtained with a 0.5 mb cross section, $a$ = 0.3, $b$ = 0.15, and $\tau$ = 30 fm/$c$ (Par1) provide a good description of the data at mid-rapidity but slightly underestimate the $v_{2}$ at forward and backward rapidity. On the other hand, the Par2 calculations with a larger cross section ($\sigma$ = 3 mb) describe the data at forward and backward rapidity but significantly overestimate the $v_{2}$ at mid-rapidity. The calculations without hadron rescatterings (Par6) are lower compared to the Par1 results, especially at mid-rapidity, while the deviation is negligible at forward and backward rapidity. Similar behavior is observed in the results obtained with larger Lund $a$ and $b$ (Par4). Furthermore, the results obtained from Par5 are significantly enhanced for all rapidity bins, indicating the impact of resonance decay on the integrated-$v_{2}$. One can see that the results from all configurations cannot simultaneously describe the data at mid- and forward/backward-rapidity. This suggests that the rapidity dependence of parameters (e.g., the initial conditions, the strength of parton interactions) needs to be considered in further developments of the model [75].
@@
Figure 8: (Color online) The comparison of the $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$, as a function of $\eta^{a}$ for 3.0 $<\eta^{b}<$ 4.0, in p–Pb collisions at 5.02 TeV between before and after nonflow subtraction. The results in 0–20% (left), 20–40% (middle) and 40–60% (right) centrality classes are all shown.
@@
As defined in Eq. 9, the $\eta$-dependent $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$ in p–Pb collisions at 5.02 TeV for 3 $<\eta^{b}<$ 4 and $5.04.4<\eta^{b}<$ 5.0 are shown in Fig. 7. The results obtained from the AMPT with all configurations are compared to the measurements from the CMS collaboration [42]. In order to compare with the data directly, we don’t apply the nonflow subtraction here. All calculations show a larger breakdown of factorization as $\eta^{a}$ increases, as observed in data. Similar to the findings in the $p_{\mathrm{T}}$-dependent flow vector decorrelation, a significant breakdown of factorization is observed when the parton scattering cross section is set to 0 (Par3), but the difference between the results of cross section $\sigma=0.5$ mb (Par1) and $\sigma=3$ mb (Par2) is small. The decorrelations obtained without hadronic scatterings (Par6) are consistent with Par1 calculations, suggesting that $\eta$-dependent decorrelation has no sensitivity to the hadronic scatterings. No significant effect of the resonance decay is observed for $\eta^{a}<$ 1, while a slight enhancement is shown for $\eta^{a}>$ 1 (Par1 vs Par5). On the other hand, varying the Lund parameters to $a=0.5$, $b=0.9$ (Par4) leads to a smaller decorrelation for both 3 $<\eta^{b}<$ 4 and 4.4 $<\eta^{b}<$ 5 (Par4 vs Par1), reflecting the effects from initial conditions. In general, the model calculations with all tunings, except for $\sigma=0$, provide a fair description of the data for $5.04.4<\eta^{b}<$ 5.0, but systematically overestimate the deviation from unity of decorrelations for 3 $<\eta^{b}<$ 4 and $\eta^{a}>$ 1.
@@
Figure 8 shows the $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$ as a function of $\eta^{a}$ in 0–20%, 20–40% and 40–60% centrality classes before and after the nonflow subtraction. The template fit method described in Eq. 10 is applied for the calculations of Eq. 8 and Eq. 9, where the contribution from the long-range jet correlations is suppressed. After subtraction, the decorrelation decreases significantly across all three centrality classes, approaching unity. This indicates that the $\eta$-dependent flow vector decorrelation in p–Pb collisions is strongly influenced by nonflow effects (e.g., dijets). This is consistent with the findings in recent measurements in smaller pp collisions performed by the ATLAS collaboration [57]. On the other hand, the difference between the results before and after the nonflow subtraction becomes larger with the increasing centrality. Especially in the 40–60% centrality class, the tendency of the decorrelation completely changes after the nonflow subtraction, first increasing then decreasing with the increasing $\eta^{a}$. It is worth noting that such an effect from the long-range jet correlations is negligible for the $p_{\mathrm{T}}$-dependent decorrelation at mid-rapidity ($|\eta|<$0.8), as shown in Fig. 1. This is because the contribution from long-range jet correlations to the $v_{2}$ is more significant at forward/backward rapidity compared to mid-rapidity in small collision systems [83]. This hints at the important role of the subtraction of long-range jet correlations in the calculations of longitudinal fluctuations of the flow vector in the small collision systems. More measurements of the flow vector decorrelation in p–Pb collisions with a proper long-range jet contribution subtration are needed in the future.

### v2

@@ IV.1 Transverse-momentum dependence of decorrelation
We first examine the $p_{\mathrm{T}}$-differential $v_{2}$ of charged hadrons in the pseudorapidity region $|\eta|<$ 0.8 using the AMPT model with the default configuration (Par1) as described in Table 1. The $v_{2}\{2\}$ and $v_{2}[2]$ are calculated according to Eq. 5 and Eq. 6. Figure 1 shows the $p_{\mathrm{T}}$-differential $v_{2}\{2\}$ and $v_{2}[2\}$ for 0–20%, 20–40%, and 40–60% centrality bins obtained from the AMPT calculations, compared with ALICE measurements. To suppress contributions from short-range jet correlations, we apply $|\Delta\eta|>0.8$. The calculations reasonably reproduce the $v_{2}$ data from central to semi-central collisions, and the ratio $v_{2}\{2\}/v_{2}[2\}$ shows deviations from unity above $\it{p}_{\mathrm{T}}\simeq$ 2 GeV/$c$. These deviations increase with $\it{p}_{\mathrm{T}}$, similar to findings in Pb–Pb collisions, indicating $p_{\mathrm{T}}$-dependent flow vector fluctuations in p–Pb collisions from the AMPT model. Unlike Pb–Pb collisions, these deviations do not significantly depend on centrality selections. To determine if these deviations are caused by non-flow effects from long-range jet correlations, we apply the template fit subtraction method described in Eq. 13. As expected, the $v_{2}$ after subtraction, shown in Fig. 1 (dashed line), is lower than the results without subtraction, especially at high $p_{\mathrm{T}}>$ 2 GeV/$c$. The ratio $v_{2}\{2\}/v_{2}[2\}$ after subtraction is consistent with that before subtraction within uncertainties. This indicates that the long-range jet contribution to the decorrelations is negligible here, and the observed $p_{\mathrm{T}}$-dependent flow vector fluctuations in p–Pb collisions are mainly from real flow signals.
@@
Figure 4: (Color online) The $v_{2}$ as a function of pseudorapidity $\eta$ in in different centrality ranges, obtained from the AMPT model with various configurations. ALICE data points are shown as black open circles for comparison.
The effect from varying the initial conditions via the Lund string parameters, $a$ and $b$, can be seen by comparing the results from Par4 and Par1, as shown in Fig. 2 (top). One can see that the $v_{2}\{2\}$$v_{2}[2]$ ratio obtained with larger Lund $a$ and $b$ (Par4) is lower in the most 0–20% collisions. With the increainsg of the centralities, the resuts from Par4 are enhanced, and finally are hihger than the results from Par1. The effects from the hadronic scatterings are studied by varying the cutoff time of hadronic rescatterings, which can seen in the comparison of the results of Par1, Par5 and Par6, as shown in Fig. 2 (bottom). The $v_{2}\{2\}$$v_{2}[2]$ ratio from Par1 and Par6 are consistent within uncertainties from the central to peripheral collisions, indicating that the impact from hadronic scatterings on the flow vector fluctuations are negligible. It is consistent with the previous findings that the hadronic scatterings have almost no effects on the $p_{\mathrm{T}}$-differential $v_{2}$ in p–Pb collisions [49]. For the case of Par5, where both the resonance decay and the hadronic scatterings are turn off, the $v_{2}\{2\}$$v_{2}[2]$ ratio is lower at $p_{\mathrm{T}}<0.8$ GeV/$c$ but higher at $p_{\mathrm{T}}>1.4$ GeV/$c$ compare to the results of Par1 and Par6. It is expected since the decay processes convert high-$p_{\mathrm{T}}$ hadrons into low-$p_{\mathrm{T}}$ particles. One can see that the results of all configurations underestimate the data, suggesting that the model needs further tuning, particularly of the initial conditions and parton scattering processes. The effect from varying the initial conditions via the Lund string parameters, $a$ and $b$, can be seen by comparing the results from Par4 and Par1, as shown in Fig. 2 (top). One can see that the $v_{2}\{2\}$$v_{2}[2]$ ratio obtained with larger Lund $a$ and $b$ (Par4) is lower in the most 0–20% collisions. With the increainsg of the centralities, the resuts from Par4 are enhanced, and finally are hihger than the results from Par1. The effects from the hadronic scatterings are studied by varying the cutoff time of hadronic rescatterings, which can seen in the comparison of the results of Par1, Par5 and Par6, as shown in Fig. 2 (bottom). The $v_{2}\{2\}$$v_{2}[2]$ ratio from Par1 and Par6 are consistent within uncertainties from the central to peripheral collisions, indicating that the impact from hadronic scatterings on the flow vector fluctuations are negligible. It is consistent with the previous findings that the hadronic scatterings have almost no effects on the $p_{\mathrm{T}}$-differential $v_{2}$ in p–Pb collisions. For the case of Par5, where both the resonance decay and the hadronic scatterings are turn off, the $v_{2}\{2\}$$v_{2}[2]$ ratio is lower at $p_{\mathrm{T}}<0.8$ GeV/$c$ but higher at $p_{\mathrm{T}}>1.4$ GeV/$c$ compare to the results of Par1 and Par6. It is expected since the decay processes convert high-$p_{\mathrm{T}}$ hadrons into low-$p_{\mathrm{T}}$ particles. One can see that the results of all configurations underestimate the data, suggesting that the model needs further tuning, particularly of the initial conditions and parton scattering processes.
The factorization ratio $r_{2}$, defined in Eq. 8, as a function of $p^{a}_{\mathrm{T}}$ in different centrality and $p^{b}_{\mathrm{T}}$ ranges, is shown in Fig. 3. The comparison between the model calculations from different AMPT configurations and the ALICE data is also shown. The model with a 0.5 mb cross section, $a$ = 0.3, $b$ = 0.15 and $\tau$ = 30 fm/$c$ (Par1), provides a fair description of the data in all centrality and $p_{\mathrm{T}}$ bins. The deviation of the $r_{2}$ from the unity has no dependence on the centrality selections, and is more pronounced with the increasing $|p^{a}_{\mathrm{T}}-p^{b}_{\mathrm{T}}|$. The $r_{2}$ obtained with larger Lund $a$ and $b$ (Par4) are higher in more peripheral collisions, reflecting the effects from the initial conditions. Changing the partonic scattering cross section from 0.5 mb to 3 mb (Par1 vs Par3) slightly enhance the $r_{2}$, while the results obtained with $\sigma=0$ mb shows a significant breakdown of the factorization at large $|p^{a}_{\mathrm{T}}-p^{b}_{\mathrm{T}}|$. The impacts from the hadronic scatterings are negligible for all centralities and $p_{\mathrm{T}}$ bins (Par1 vs Par6), but the resonance decay process still has large impacts on the $r_{2}$ at low-$p_{\mathrm{T}}$ region (Par1 vs Par5). One can see that the calculations of $r_{2}$ and $v_{2}\{2\}$$v_{2}[2]$ ratio have generally similar dependence on the initial conditions, parton scattering cross section, hadron scattering process. It is expected since $r_{2}$ is basically the double-differential $v_{2}\{2\}$$v_{2}[2]$ ratio.
@@
Figure 5: (Color online) The square root of the product of factorization ratios, $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$, as a function of $\eta^{a}$ for 3.0 $\eta^{b}<$ 4.0 (left) and $4.4<\eta^{b}<5.0$ (right), in 0–20% centrality class of p–Pb collisions at 5.02 TeV, obtained from the AMPT model with various tunings. The CMS data points are shown as black open circles for comparison.
@@ IV.2 Pseudorapidity-dependence of decorrelation
The $\eta$-dependent $v_{2}$ is first calculated before studying the decorrelation. We follow the two-particle correlation method used by the ALICE experiment. The correlation between three groups of particles is constructed based on Eq. 13, then the flow coefficients are extracted for three combinations:
$aligned v_{n}(\eta^{a})=\sqrt{\frac{V_{n\Delta}(\eta^{a},\eta^{b})V_{n\Delta}(\eta^{a},\eta^{c})}{V_{n\Delta}(\eta^{b},\eta^{c})}},$ (14)
@@
where $a$, $b$, and $c$ represent the hadrons with $p_{\mathrm{T}}>$ 0 GeV/$c$ at different rapidity bins. Similarly, $v_{n}(\eta^{a})$ and $v_{n}(\eta^{b})$ can be obtained by alternating the indices $a$, $b$, and $c$ in Eq. 14. Figure 4 shows the $v_{2}$ of the charged hadrons for $p_{\mathrm{T}}>$ 0 GeV/$c$ as a function of $\eta$ in 0–5%, 5–10%, 10–20%, and 20–40% centrality ranges. The results obtained from different configurations of the AMPT model are compared with the ALICE data. The calculations obtained with a 0.5 mb cross section, $a$ = 0.3, $b$ = 0.15, and $\tau$ = 30 fm/$c$ (Par1) provide a good description of the data at mid-rapidity but slightly underestimate the $v_{2}$ at forward and backward rapidity. On the other hand, the Par2 calculations with a larger cross section ($\sigma$ = 3 mb) describe the data at forward and backward rapidity but significantly overestimate the $v_{2}$ at mid-rapidity. The calculations without hadron rescatterings (Par6) are lower compared to the Par1 results, especially at mid-rapidity, while the deviation is negligible at forward and backward rapidity. Similar behavior is observed in the results obtained with larger Lund $a$ and $b$ (Par4). Furthermore, the results obtained from Par5 are significantly enhanced for all rapidity bins, indicating the impact of resonance decay on the integrated-$v_{2}$. One can see that the results from all configurations cannot simultaneously describe the data at mid- and forward/backward-rapidity. This suggests that the rapidity dependence of parameters (e.g., the initial conditions, the strength of parton interactions) needs to be considered in further developments of the model [39].
@@
Figure 6: (Color online) The comparison of the $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$, as a function of $\eta^{a}$ for 3.0 $\eta^{b}<$ 4.0, in p–Pb collisions at 5.02 TeV between before and after nonflow subtraction. The results in 0–20% (left), 20–40% (middle) and 40–60% (right) centrality classes are all shown.
@@
As defined in Eq. 12, the $\eta$-dependent $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$ in p–Pb collisions at 5.02 TeV for 3 $\eta^{b}<$ 4 and $4.4<\eta^{b}<5.0$ are shown in Fig. 5. The results obtained from the AMPT with all configurations are compared to the measurements from the CMS collaboration. All calculations show a larger breakdown of factorization as $\eta^{a}$ increases, consistent with the data. Similar to the findings in the $p_{\mathrm{T}}$-dependent flow vector decorrelation, a significant breakdown of factorization is observed when the parton scattering cross section is set to 0 (Par3), but the deviation between the results of cross section $\sigma=0.5$ mb (Par1) and $\sigma=3$ mb (Par2) is small. Unlike the integrated-$v_{2}$ in Fig. 4, the decorrelations obtained without hadronic scatterings (Par6) are consistent with Par1 calculations, suggesting that $\eta$-dependent decorrelation has no sensitivity to the hadronic scatterings. No significant effect of the resonance decay is observed for $\eta^{a}<$ 1, while a slight enhancement is shown for $\eta^{a}>$ 1 (Par1 vs Par5). On the other hand, varying the Lund parameters to $a=0.5$, $b=0.9$ (Par4) leads to a smaller decorrelation for both 3 $\eta^{b}<$ 4 and 4.4 $\eta^{b}<$ 5 (Par4 vs Par1), reflecting the effects from initial conditions. In general, the model calculations with all tunings, except for $\sigma=0$, provide a fair description of the data for $4.4<\eta^{b}<5.0$, but systematically overestimate the deviation from unity of decorrelations for 3 $\eta^{b}<$ 4 and $\eta^{a}>$ 1.
@@
Figure 6 shows the $\sqrt{r_{2}(\eta^{a},\eta^{b})r_{2}(-\eta^{a},-\eta^{b})}$ as a function of $\eta^{a}$ in 0–20%, 20–40% and 40–60% centrality classes before and after the nonflow subtraction. The template fit method described in Eq. 13 is applied for the calculations of Eq. 10 and Eq. 12, where the contribution from the long-range jet correlations is suppressed. A significant enhancement for the decorrelation is observed for the results after the subtraction in three centrality classes, indicating that the $\eta$-dependent flow vector decorrelation has a strong dependence on the nonflow (e.g., the dijet) in p–Pb collisions. This is consistent with the findings in recent measurements in even smaller pp collisions performed by the ATLAS collaboration [1]. On the other hand, the deviation between the results before and after the nonflow subtraction becomes larger with the increasing centrality. Especially in the 40–60% centrality class, the tendency of the decorrelation completely changes after the nonflow subtraction, first increasing then decreasing with the increasing $\eta^{a}$. It is worth noting that such an effect from the long-range jet correlations is negligible for the $p_{\mathrm{T}}$-dependent decorrelation at mid-rapidity ($|\eta|<$0.8), as shown in Fig. 1. This is because the contribution from long-range jet correlations to the $v_{2}$ is more significant at forward/backward rapidity compared to mid-rapidity in small collision systems [12]. This hints at the important role of the long-range jet correlations in the longitudinal fluctuations and decorrelations of the flow vector.